Two sides if a triangle have lengths 4cm and 5 cm. The 3rd side has x cm, where x is a positive integer. How many different values of x are there?
I know that you can use pythagoras for this ut I'm not sure how as you aren't told which side is which.

Answer by rothauserc(4718)

this is a 3, 4, 5 right triangle
3^2 + 4^2 = 5^2
9 +16 = 25
x = 3 cm